OpenStudy (anonymous):

Find f '(x) and f ''(x). f(x) = x^8ex f '(x) = f ''(x) =

OpenStudy (anonymous):

\[f(x)=x^8e^x\]?

OpenStudy (anonymous):

f(x) = x^8e^x

OpenStudy (goformit100):

Diffrentiate it and reply your answer

OpenStudy (anonymous):

product rule for this one use \[(fg)'=f'g+g'f\] with \(f(x)=x^8,f'(x)=8x^7,g(x)=e^x,g'(x)=e^x\)

OpenStudy (anonymous):

(8x^7)(e^x)+(x^8)(e^x)

OpenStudy (cruffo):

that answer is good for f'(x).

OpenStudy (anonymous):

how do i do f''(x)

OpenStudy (cruffo):

\(\large f'(x) = 8x^7e^x + x^8e^x\) just take the derivative of each term in the sum above using product rule

OpenStudy (anonymous):

f''(x) = (56x^6)(e^x)+(8x^7)(e^x)

OpenStudy (cruffo):

that is the derivative of the first term, but not f''(x) yet.

OpenStudy (anonymous):

f''(x) = ((56x^6)(e^x)+(8x^7)(e^x))+((8x^7)(e^x)+(8x^7)(e^x))

OpenStudy (cruffo):

good except for the last term in you answer \[f''(x) = 56^6e^x + 8x^7e^x + 8x^7e^x + x^8e^x\]

OpenStudy (cruffo):

sorry, missed an x :) \[f''(x) = 56x^6e^x + 8x^7e^x + 8x^7e^x + x^8e^x\]

OpenStudy (anonymous):

thank you

OpenStudy (cruffo):

you may need to gather like terms...

OpenStudy (anonymous):

okaay

OpenStudy (anonymous):

would it be 16x^7(e^x)

OpenStudy (cruffo):

yep!

OpenStudy (anonymous):

i got it wrong

OpenStudy (cruffo):

maybe you also need to factor it??? \[f''(x) = 56x^6e^x + 16x^7e^x + x^8e^x\] \[GCF = x^6e^x\]

OpenStudy (anonymous):

(56x^6)(e^x)+(x^8)(e^x)+(16x^7)(e^x)

OpenStudy (cruffo):

maybe you also need to factor it??? \[f''(x) = 56x^6e^x + 16x^7e^x + x^8e^x\] \[GCF = x^6e^x\]

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

have to simplify

OpenStudy (anonymous):

i got ti thank you

OpenStudy (cruffo):

cool.